OpenStudy (anonymous):

Divide using synthetic division. -2x^3+11x^2-x+20 by (x-4)

Nnesha (nnesha):

for synthetic division you have to solve x-4 for x do it x-4 = 0 solve for x

Nnesha (nnesha):

 just write COEFFICIENT and constant term forget about variables for a sec

Nnesha (nnesha):

x-4 = 0 x= ??? :-)

OpenStudy (anonymous):

Thank you very much! ;)

Nnesha (nnesha):

hmm we are not done yet ;-) do you know how to solve x-4 for x ?

Nnesha (nnesha):

feel free to say no ;-)

OpenStudy (anonymous):

Add 4 to both sides

Nnesha (nnesha):

great :-)

Nnesha (nnesha):

x- 4 = 0 add 4 both side you will get x =4 now we have to multiply all coefficient by 4  carry down first term